Given that the standard concentration formula is (M1V1=M2V2) and the Cu initial concentration is 63.55 mg/mL plus data table info:

What would be V1 and V2? Please show 2 examples from data table info.

Solution Starting buret read. Ending buret read. Total Delivered concentration
1 mL 22.89 mL 23.91 mL 1.02 mL ?
2 mL 15.51 mL 17.49 mL 1.98 mL ?
4 mL 17.49 mL 21.69 mL 4.20 mL ?
6 mL 11.91 mL 17.92 mL 6.01 mL ?
8 mL 13.59 mL 21.56 mL 7.97 mL ?
Answer #1

Lets consider the 1st and the 4th readings.

In the first reading, M1 = 63.55 mg/ml, V1= 1.02 ml, V2 = 1ml

M1*V1 = M2*V2

63.55*1.02 = M2*1

M2 = 64.821 mg/ml

In the fourth reading, M1=63.55mg/ml, V1= 4.2ml, V2 = 4ml

M1*V1 = M2*V2

63.55 * 4.2 = M2 * 4

M2 = 66.727 mg/ml